OpenStudy (anonymous):

A cyclist maintains a constant velocity of 4.3 m/s headed away from point A. At some initial time, the cyclist is 255 m from point A. What will be his displacement from his starting position after 53 s? Answer in units of m What will be his position from point A after that time? Answer in units of m

OpenStudy (john_es):

Suppose the initial position is x0=255 m, then, \[\Delta x=(x_0+vt)-x_0=4.3\cdot53=227.9\ m\] From point A, \[x=x_0+vt=255+227.9=482.9\ m\]

OpenStudy (anonymous):

I got the same answer but it isn't correct!

OpenStudy (john_es):

Both answers?

OpenStudy (john_es):

Well, lets suppose the cyclist is going to point A, and at some initial time, he is 255 m away from A. In other words, his positions is x=-255 m. After 53 s, the position of the cyclist will be, \[x=x_0+vt=-255+4.3\cdot53=-27.1\ m\] so his position now is 27.1 m away from point A (he still has not reached point A). However the displacement will be, \[\Delta x=-16.5-(-255)=227.9\ m\]